# FALCONEye: Finding Answers and Localizing Content in ONE-hour-long videos with multi-modal LLMs

### v1

@@ 3 FALCONEye
We introduce FALCONEye, a meta-architecture that combines a VLM and an LLM within an exploration-based search algorithm to tackle the VAS task. Unlike previous agents designed for MCQs and short videos, operating at the frame level, FALCONEye is built to handle OQs and one-hour-long videos, operating at the clip level.
Figure 2: FALCONEye exploration algorithm that follows our FALCONEye meta-architecture. Given a question ( Refer to caption) and a video, it starts with a global Pre-processing of the video, where the VLM generates short captions ( Refer to caption) from video clips. Afterwards, the LLM Reasons from the Q and the captions to select some candidate clips to explore. The VLM Evaluates the candidate clips sampling frames varying number and resolution according to the clip duration. We get an answer ( Refer to caption) and its confidence ( Refer to caption) for each candidate clip. From all { Refer to caption, Refer to caption, Refer to caption} tuples, the LLM Decides wether the final answer has already been found or, otherwise, it returns some promising clips to continue exploring. If these promising clips are large enough, we segment them to generate new shorter candidate clips to evaluate. If the promising clips are already too short ( Refer to caption), we ask again the LLM to reason removing the caption ( Refer to caption) of the already explored clip.
Problem definition:
Given a video-question pair ($\mathcal{V}$, $Q$). The goal of the VAS task is:
@@ 3.1 Meta-architecture
We propose a meta-architecture with two main components: a Vision-Language Model (VLM) and a Large Language Model (LLM). This meta-architecture is agnostic to any specific LLM or VLM, allowing flexibility in model selection, including frozen foundation models. Our overall idea involves using a lightweight VLM to process the most computationally expensive data (visual information) and a medium-sized LLM to handle text-based reasoning, which is significantly cheaper to process. This setup enables higher efficiency and cost reduction compared to addressing the task with current large-scale state-of-the-art VLMs. The VLM is used only to process short video clips, namely for two tasks, provide captions and generate potential answers to $Q$. The LLM powerful reasoning capabilities are exploited to extract information from all the captions and find interesting clips, and decide wether to explore other clips, refine the current clip search (zooming into it) or if the final answer has been found.
@@ 3.2 Exploration Algorithm
Evaluation-. In this stage, each candidate clip $v^{*}_{i}$ is evaluated by the VLM. First, a uniform sampling of frames is performed within the clip. The number and resolution of sampled frames vary according to the duration of $v^{*}_{i}$, creating a zoom-in effect. For first-level clips, a larger number of frames with lower resolution are sampled, whereas for shorter clips, fewer frames with higher resolution are selected. For fixed input resolution VLMs, we use tiling to increase the resolution (similar to what GPT4o [32] does).
The VLM takes as input the question $Q$ along with the sampled frames from $v^{*}_{i}$ and generates a potential answer $\mathcal{A}^{*}_{i}$ and a confidence score $p(\mathcal{A}^{*}_{i})$. The answer is derived from a set of generated tokens $\mathcal{T}\in D$, such that $\mathcal{A}^{*}_{i}=Tokenizer(\mathcal{T})$, where $D=\{t_{1},\ldots,t_{d}\}$ is the token dictionary used by the LLM. During the VLM generation process, we store the logits $\{l_{1}^{k},\ldots,l_{d}^{k}\}$ for each generated token $t_{k}\in\mathcal{T}$ and compute $p(\mathcal{A}^{*}_{i})$ as the geometric average of the answer token probabilities [26]:
$p(\mathcal{A}^{*}_{i})=\sqrt[s]{\prod_{k=1}^{s}\max\left(Softmax\left(\{l_{1}^{k},\ldots,l_{d}^{k}\}\right)\right)},$ (1)
where $s$ is the number of generated tokens in $\mathcal{T}$ (see Fig. 9 in supp. material). In MCQ scenario, there is no need for an average as $s=1$ (a single token for the option letter: A, B, C, D, …) and the confidence is just the softmax of the logits for the option letter token [10]. This step is supported by our VLM calibration analysis presented later in Section 6.
@@ 4 FALCON-Bench
The proposed benchmark is specifically designed for the VAS task, using one-hour-long videos obtained from external public datasets. Our contribution lies in the constructed set of Q&A that assess the VAS task on these videos. Each question is carefully crafted to ensure a unique answer within the video, with the answer content embedded in a single, short temporal window. Unlike previous VQA benchmarks, which often rely on broad video-level reasoning, we highlight the challenge of finding a needle in a haystack within long-form videos. Our benchmark emphasizes the necessity of not just uniformly sampling frames or applying token compression across the entire video, but instead reasoning over the content to localize the exact clip that contains the answer. To facilitate the evaluation of the localization subtask, we provide the ground-truth (GT) temporal interval where the answer is contained.
The videos of our benchmark were sourced from SoccerNet [16], MovieChat-1K [38] and Walking Tours Dataset [40], ensuring diversity and comprehensive coverage across different video types. Overall, the benchmark comprises 575 questions, covering 4 categories, over 90 videos, with an average video duration of 78.9 minutes and answers localized within a GT temporal window of 38.4 seconds. The dataset is split into a test set (506 questions) and a validation set (70 questions). Each question is accompanied by four different answer options, allowing for multiple-choice question (MCQs) evaluation. However, the primary focus remains on open-ended questions (OQs), as this approach is more aligned with the VAS task. The annotation process was conducted by humans assisted by GPT-4o for generating answer options. More details on the benchmark content and construction process are in Sec. 8 of the supplementary material. The benchmark has been built under the lmms-eval framework [51] to ease further evaluation and will be publicly available upon acceptance.
@@
Open-Source VLMs with Multi-Image Support
mPLUG-Owl3 [47] 7B 16 5.52 41.1 28.5 36.0 35.2 N/A 6.63 8.82 4.23 10.0 7.68 0.68 0.29 0.66 0.54 N/A
LLaVA-v1.5 [25] 7B 32 11.4 23.5 31.9 44.0 33.1 N/A 12.7 3.92 4.80 2.00 3.57 0.36 0.31 0.24 0.30 N/A
LLaVA-v1.6 [19] 7B 32 12.9 28.4 30.5 38.0 32.3 N/A 14.3 4.90 4.51 2.00 3.80 0.31 0.30 0.20 0.27 N/A
LLaVA-OV [17] 7B 32 16.9 48.0 30.2 46.0 41.4 N/A 7.29 9.80 4.23 16.0 10.0 0.53 0.25 1.00 0.59 N/A
Qwen2.5-VL [4] 7B 768 68.9 50.0 24.4 48.0 40.8 N/A 70.1 11.7 8.19 10.0 9.96 0.72 0.53 0.62 0.62 N/A
Open-Source VLMs designed for videos
Video-LLaVA [22] 7B 8 3.37 28.4 26.8 34.0 29.7 N/A 4.50 6.86 5.08 8.00 6.64 0.56 0.37 0.68 0.53 N/A
VideoChat2-HD [21] 7B 8 6.74 29.4 24.2 26.0 26.5 N/A 8.06 8.82 2.82 8.00 6.54 0.71 0.20 0.52 0.47 N/A
LLaVA-Video [52] 7B 32 12.4 50.0 38.9 58.0 48.9 N/A 13.0 10.7 11.0 12.0 11.2 0.76 0.66 0.86 0.76 N/A
Open-Source VLMs specific for long videos
MovieChat-OV [38] 7B 1024 62.3 43.1 38.7 46.0 42.6 N/A 69.3 9.80 4.80 10.0 8.20 0.53 0.25 0.66 0.48 N/A
Apollo [54] 7B 0.5fps 36.2 51.9 35.8 64.0 50.5 N/A 36.5 16.6 10.7 18.0 15.1 0.96 0.68 1.06 0.90 N/A
Meta-architectures built from a LLM (GPT4o-mini) and a VLM (Qwen2.5-VL)
Socratic [50] 7B+N/A 0.5fps 186.6 53.9 48.0 54.0 51.9 21.3 186.7 25.4 29.0 20.0 24.8 1.49 1.71 1.16 1.45 19.6
VideoAgent++ [43] 7B+N/A 9.2 45.0 48.0 52.2 36.0 45.4 19.1 43.8 8.82 24.8 8.00 13.8 0.56 1.38 0.44 0.79 13.3
FALCONEye [Uncaptioned image] 7B+N/A 659.2 393.0 74.5 61.5 74.0 70.0 27.7 245.4 50.9 37.2 46.0 44.7 2.67 1.96 2.50 2.38 24.9
FALCONEye [Uncaptioned image] 7B+N/A 518.5 213.3 67.6 58.4 68.0 64.7 25.2 187.9 43.1 38.4 42.0 41.1 2.36 2.03 2.28 2.22 22.7
@@ 5 Experiments
In MCQs, we report accuracy as the primary evaluation metric. For OQs, following the approach of Maaz et al. [27], also applied in others, as InfiniBench [3], we employ GPT-assisted evaluation. GPT reports accuracy (100 or 0) and a score value (0-5 range), assessing key aspects such as Correctness of Information, Detailed Orientation, Contextual Understanding, Temporal Understanding and Consistency. We also use the Ground Truth over Union (GToU) metric to evaluate the quality of the predicted temporal interval, as a visual evidence (short clip) that contains the answer. Section 8.3 of the Supplementary discusses why we find GToU more adequate than mIoU (a clip that tries to match the entire GT clip).
Table 3: Meta-architectures cost and performance comparison in the test split of the FALCON-Bench OQs. We set LLM=GPT-4o-mini, VLM=Qwen2.5-VL. We measure LLM usage as number of inferences (#Inf.) per question and, total tokens (#Toks) and time (s) per inference. We measure VLM usage as number of inferences (#Inf.T) per question, and number of frames (#F), max output tokens (#T) and time (s) per inference. We compute #Inf.T adding the #Inf per Q and per V as #$\#Inf.T=\#Inf.Q+\#Inf.V/Q.V$, knowing that $Q.V=6.24$. VLM inferences can be for Captioning or VQA. We report the average accuracy (Acc.), score (Sc.) and localization (Loc.).
LLM VLM for Captioning or VQA Total Performance
Usage Avg. per Inf #Inf. Avg. per Inf Time Open Questions
Model Name Mode #Inf. #Toks s V Q T #F #T s s Acc. Sc. Loc.
Socratic [50] QA 1 13.5K 2.1 78.9 - 12.6 30 768 14.6 186.7 24.8 1.45 19.6
Socratic [Uncaptioned image] [50] QA 1 2.4K 1.1 78.9 - 12.6 30 64 9.8 125.0 15.5 1.03 18.3
VideoAgent++ [43] QA 5.6 1.3K 3.2 5 2.90 3.70 1 64 2.2 43.8 13.8 0.79 13.3
SequentialBP - - - - - 53.3 53.3 120 64 12.7 676.9 28.7 1.52 13.7
Sequential - - - - - 78.9 78.9 120 64 12.7 1002 34.8 1.83 9.05
IterativeSamp - - - - - 85.4 85.4 120 64 18.7 1596 24.0 1.34 1.26
FALCONEye [Uncaptioned image] Reason 4.3 1.7K 5.8 78.9 9.98 22.5 29.3 64 9.8 245.4 44.7 2.50 24.9
FALCONEye [Uncaptioned image] Reason 3.0 1.7K 5.8 78.9 4.86 17.4 29.8 64 9.8 187.9 41.8 2.24 22.7
Illustration of exploration baselines
Exploration Baselines-.
A widely used baseline in recent long VQA benchmarks [5] is the Socratic method [50]. In this baseline, a VLM generates captions $\mathcal{C}=\{c_{1},\ldots,c_{n}\}$ from short video clips $\{v_{1},\ldots,v_{n}\}\in\mathcal{V}$ and a LLM generates $\mathcal{A}$ from $Q$ and $\mathcal{C}$. In addition to this baseline, we introduce three alternative exploration baselines (Fig. within Tab. 3) that rely solely on a VLM, guided by its answer confidence computed as described in Sec 3.
@@
We evaluate our meta-architecture using Qwen2.5-VL [4] as the VLM and GPT4o-mini [31] as the LLM. As detailed in Sec. 6, Qwen2.5-VL outperforms other small-size VLMs on FALCONBench OQs in short clips near the answer, and is well calibrated. Its ViT supports images of any resolution, dynamically adjusting visual tokens, enabling an accurate zoom-in effect as clip lengths decrease. GPT4o-mini provides a balanced trade-off between cost and reasoning power (Sec. 6). Prompts used for LLM queries are detailed in Sec. 9.2. Inspired by the Socratic approach [5], first-level clips are 1min length, but captions are now shorter formed by up to 64 tokens (instead of 768). Within these clips, second-order (20s) and third-order (5s) clips are explored. The zoom-in effect, validated in Sec. 6, adjusts from 30 frames ($672\times 1204$) in first-level clips to 10 frames ($824\times 1462$) in third-order clips. In OQs, the LLM determines the final answer based on completion, confidence ($0.8$ threshold), captions, and temporal localization (as explained in Sec. 3). In MCQs, this evaluation is replaced by a simple confidence-based check with $0.9$ threshold since all possible answers are predefined (A, B, C, or D) and inherently valid, making confidence the only selection criterion. Lastly, we evaluate two versions of FALCONEye: FALCONEye [Uncaptioned image], the top-performing configuration, which evaluates up to 45 candidate clips, and FALCONEye [Uncaptioned image], a cost-efficient variant that limits evaluation to 10 candidates.
Table 4: Model performance comparison in LongVideoBench and MLVU benchmarks.
Model MLVU tasks LongVideoBench tasks *
Name S # F NQA ER PQA S2E S2O S2A E2O O2E
858s 575s 466s 727s 716s 642s 470s 672s
Baselines
Full Mark - - 100 100 100 100 100 100 100 100
Random - - 16.7 16.7 16.7 20.0 20.0 20.0 20.0 20.0
Proprietary Long-Context LLMs
Gemini-1.5-Pro N/A 256 N/A N/A N/A 74.6 58.3 76.2 68.7 73.3
GPT-4o N/A 0.5fps 42.9 47.8 57.1 76.8 69.8 70.9 67.3 72.8
Open-Source Image LLMs with Multi-Image Support
LLaVA-OV [17] 72B 32 46.7 58.4 58.0 N/A N/A N/A N/A N/A
InternVL2.5 [7] 72B 64 48.3 47.2 52.0 N/A N/A N/A N/A N/A
LLaVA-Video [52] 34B 32 N/A N/A N/A 62.1 50.2 51.2 50.9 58.5
Open-Source Video LLMs
VILA-1.5 [23] 40B 14 38.3 35.8 62.0 N/A N/A N/A N/A N/A
Aria [18] 25B 256 56.6 66.0 58.3 N/A N/A N/A N/A N/A
Meta-architectures built from a LLM (GPT4o-mini) and a VLM (Qwen2.5-VL)
FALCONEye [Uncaptioned image] 7B+N/A 2fps 61.7 43.5 62.9 60.2 43.1 64.8 60.0 64.4
FALCONEye [Uncaptioned image] 7B+N/A 2fps 62.0 40.9 59.6 59.1 38.9 63.6 53.8 62.1
– : Does not apply. N/A: Not Available due to unknown model size or not published results on benchmark.
* Our results are in validation split since gt test split is not publicly available.

### v2

@@ 3 FALCONEye
FALCONEye is a novel video agent combining a training-free and model-agnostic meta-architecture (VLM+LLM) with an exploration-based search algorithm guided by calibrated confidence from VLM’s answers. Unlike previous agents designed for MCQs, short videos and frame-level operations, FALCONEye handles OQs, one-hour-long videos, and operates at clip level.
Figure 2: FALCONEye exploration algorithm. Given a question ( Refer to caption) and a video, it starts with a global Pre-processing of the video, where the VLM generates short captions ( Refer to caption) from video clips. Afterwards, the LLM Reasons from the Q and the captions to select some candidate clips to explore. The VLM Evaluates the candidate clips sampling frames varying number and resolution according to the clip duration. We get an answer ( Refer to caption) and its confidence (Conf.) for each candidate clip. From all { Refer to caption, Conf., Refer to caption} tuples, the LLM Decides wether the final answer has already been found or to continue exploring with two different ways: if some candidate clips are still promising and large enough, we segment them to generate new shorter candidate clips to evaluate. Otherwise, we ask again the LLM to reason removing the caption of the already explored clip.
VAS Problem definition:
Given a video-question pair ($\mathcal{V}$, $Q$), the goal of the VAS task is:
@@ 3.1 Meta-architecture
We propose a meta-architecture with two main components: a VLM and an LLM. This meta-architecture is training-free and agnostic to any specific LLM or VLM, allowing flexibility in model selection. Our overall idea involves using a small VLM to process the most computationally expensive data (visual information) and a medium-sized LLM to handle text-based reasoning, which is significantly cheaper to process. This setup enables higher efficiency and cost reduction compared to addressing the task with current large-scale state-of-the-art VLMs. The VLM is used only to process short video clips, namely for two tasks, provide captions and generate potential answers to $Q$. The LLM powerful reasoning capabilities are exploited to extract information from all the captions and find interesting clips, and to decide wether to explore other clips, refine the current clip search (zooming into it) or stop the search.
@@ 3.2 Exploration Algorithm
Evaluation-. In this stage, each candidate clip $v^{*}_{i}$ is evaluated by the VLM. First, a uniform sampling of frames is performed within the clip. The number and resolution of sampled frames vary according to the duration of $v^{*}_{i}$, creating a zoom-in effect. For first-level clips, a larger number of frames with lower resolution are sampled, whereas for shorter clips, fewer frames with higher resolution are selected. For fixed input resolution VLMs, we use tiling to increase the resolution (similar to what GPT4o [openai2023gpt4v] does).
The VLM takes as input the question $Q$ along with the sampled frames from $v^{*}_{i}$ and generates a potential answer $\mathcal{A}^{*}_{i}$ and a confidence score $p(\mathcal{A}^{*}_{i})$. The answer is derived from a set of generated tokens $\mathcal{T}\in D$, such that $\mathcal{A}^{*}_{i}=Tokenizer(\mathcal{T})$, where $D=\{t_{1},\ldots,t_{d}\}$ is the token dictionary used by the LLM. During the VLM generation process, we store the logits $\{l_{1}^{k},\ldots,l_{d}^{k}\}$ for each generated token $t_{k}\in\mathcal{T}$ and compute $p(\mathcal{A}^{*}_{i})$ as the geometric average of the answer token probabilities [liu2023litcab]:
$\footnotesize p(\mathcal{A}^{*}_{i})=\sqrt[s]{\prod_{k=1}^{s}\max\left(Softmax\left(\{l_{1}^{k},\ldots,l_{d}^{k}\}\right)\right)},$ (1)
where $s$ is the number of generated tokens in $\mathcal{T}$ (see Fig. 9 in supp. material). In MCQ scenario, there is no need for an average as $s=1$ (a single token for the option letter: A, B, C, D, …) and the confidence is just the softmax of the logits for the option letter token [chung2023long]. This step is supported by our VLM calibration analysis presented later in Section 6.
@@ 4 FALCON-Bench
FALCON-Bench is the first benchmark specifically tailored to the VAS task, where models must not only provide an answer to a question, but also retrieve the precise temporal window that supports it. Our benchmark raises the difficulty across all axes: we use one-hour-long videos, we focus on open-ended questions, and evaluate both answer quality and temporal grounding jointly. While all videos were obtained from external public datasets, our contribution lies in the constructed set of Q&A that assess the VAS task on these videos, where each question is carefully crafted to ensure a unique answer within the video, with the answer content embedded in a single, short temporal window. Unlike previous VQA benchmarks, which often rely on broad video-level reasoning, we highlight the challenge of finding a needle in a haystack within long-form videos. Our benchmark emphasizes the necessity of not just uniformly sampling frames or applying token compression across the entire video, but instead reasoning over the content to localize the exact clip that contains the answer.
The videos of our benchmark were sourced from SoccerNet [giancola2018soccernet], MovieChat-1K [song2024moviechat] and Walking Tours Dataset [venkataramanan2023imagenet], ensuring diversity and comprehensive coverage across different video types. Overall, the benchmark comprises 575 questions, covering 4 categories, over 90 videos, with an average video duration of 78.9 minutes and answers localized within a GT temporal window of 38.4 seconds. The dataset is split into a test set (506 questions) and a validation set (70 questions). Each question is accompanied by four different answer options, allowing for multiple-choice question (MCQs) evaluation. However, the primary focus remains on open-ended questions (OQs), as this approach is more aligned with the VAS task. The annotation process was conducted by humans assisted by GPT-4o for generating answer options. More details on the benchmark content and construction process are in Sec. 8 of the supplementary material. The benchmark has been built under the lmms-eval framework [zhang2024lmmseval] to ease further evaluation and will be publicly available upon acceptance.
@@
[…]
mPLUG-Owl3 [ye2024mplug] 7B 16 5.52 41.1 28.5 36.0 35.2 0.00 6.63 8.82 4.23 10.0 7.68 0.68 0.29 0.66 0.54 0.00
LLaVA-v1.5 [liu2023llava] 7B 32 11.4 23.5 31.9 44.0 33.1 0.00 12.7 3.92 4.80 2.00 3.57 0.36 0.31 0.24 0.30 0.00
LLaVA-v1.6 [llava-next] 7B 32 12.9 28.4 30.5 38.0 32.3 0.00 14.3 4.90 4.51 2.00 3.80 0.31 0.30 0.20 0.27 0.00
LLaVA-OV [llava-onevision] 7B 32 16.9 48.0 30.2 46.0 41.4 0.00 7.29 9.80 4.23 16.0 10.0 0.53 0.25 1.00 0.59 0.00
Qwen2.5-VL [Qwen2.5-VL] 7B 768 68.9 50.0 24.4 48.0 40.8 0.00 70.1 11.7 8.19 10.0 9.96 0.72 0.53 0.62 0.62 0.00
[…]
Video-LLaVA [lin2023video] 7B 8 3.37 28.4 26.8 34.0 29.7 0.00 4.50 6.86 5.08 8.00 6.64 0.56 0.37 0.68 0.53 0.00
VideoChat2-HD [li2024mvbench] 7B 8 6.74 29.4 24.2 26.0 26.5 0.00 8.06 8.82 2.82 8.00 6.54 0.71 0.20 0.52 0.47 0.00
LLaVA-Video [zhang2024llavanext-video] 7B 32 12.4 50.0 38.9 58.0 48.9 0.00 13.0 10.7 11.0 12.0 11.2 0.76 0.66 0.86 0.76 0.00
[…]
MovieChat-OV [song2024moviechat] 7B 1024 62.3 43.1 38.7 46.0 42.6 0.00 69.3 9.80 4.80 10.0 8.20 0.53 0.25 0.66 0.48 0.00
Apollo [zohar2024apollo] 7B 0.5fps 36.2 51.9 35.8 64.0 50.5 0.00 36.5 16.6 10.7 18.0 15.1 0.96 0.68 1.06 0.90 0.00
Video-XL-Pro [liu2025video] 3B 256 48.8 52.9 33.8 50.0 45.6 0.00 7B 17.6 7.62 14.0 13.0 1.08 0.45 0.94 0.63 0.00
LLaVA-OV+ReKV [di2025streaming] 7B 0.5fps 92.1 65.6 40.1 56.0 53.9 0.00 96.6 34.3 13.8 24.0 24.0 1.78 0.80 1.40 1.32 0.00
[…]
Socratic [zeng2022socratic] 7B+N/A 0.5fps 186.6 53.9 48.0 54.0 51.9 21.3 186.7 25.4 29.0 20.0 24.8 1.49 1.71 1.16 1.45 19.6
LifelongMemory[wang2024lifelongmemory] 7B+N/A 0.5fps 207.3 45.0 41.5 38.0 41.8 0.00 208.2 6.86 22.3 8.00 12.3 0.53 1.37 0.56 0.82 0.00
VideoAgent[wang2024videoagent]+OQ 7B+N/A 9.2 45.0 48.0 52.2 36.0 45.4 19.1 43.8 8.82 24.8 8.00 13.8 0.56 1.38 0.44 0.79 13.3
FALCONEye-Pro 7B+N/A 659.2 393.0 74.5 61.5 74.0 70.0 27.7 245.4 50.9 37.2 46.0 44.7 2.67 1.96 2.50 2.38 24.9
FALCONEye-Flash 7B+N/A 518.5 213.3 67.6 58.4 68.0 64.7 25.2 187.9 43.1 38.4 42.0 41.1 2.36 2.03 2.28 2.22 22.7
@@ 5 Experiments
In MCQs, we report accuracy as the primary evaluation metric. For OQs, following [maaz2023video], also applied in others, as InfiniBench [ataallah2024infinibench], we employ GPT-assisted evaluation. GPT reports accuracy (100 or 0) and a score value (0-5 range), assessing key aspects (Correctness of Information, Detailed Orientation, Contextual Understanding, Temporal Understanding and Consistency). We also use Ground Truth over Union (GToU) metric to evaluate the quality of the predicted temporal interval, as visual evidence (short clip) that contains the answer. Section 8.3 of the Supplementary discusses why we find GToU more adequate than mIoU (a clip that tries to match the entire GT clip).
Table 3: Meta-architectures cost and performance comparison in the test split of the FALCON-Bench OQs. We set LLM=GPT-4o-mini, VLM=Qwen2.5-VL. We measure LLM/VLM usage as number of inferences (#Inf.) per question and, total tokens (#Toks) and time (s) per inference. We report the average accuracy (Acc.), score (Sc.) and localization (Loc.).
LLM VLM Total Open Questions
Model Name #Inf. #Toks s #Inf. #T s s Acc. Sc. Loc.
Socratic [zeng2022socratic] 1 13.5K 2.1 12.6 768 14.6 186.7 24.8 1.45 19.6
Socratic-Short [zeng2022socratic] 1 2.4K 1.1 12.6 64 9.8 125.0 15.5 1.03 18.3
LifelongMemory[wang2024lifelongmemory] 10.2 0.3K 1.0 12.6 64 9.8 208.2 12.3 0.82 0.00
VideoAgent[wang2024videoagent]+OQ 5.6 1.3K 3.2 3.70 64 2.2 43.8 13.8 0.79 13.3
SequentialBP - - - 53.3 64 12.7 676.9 28.7 1.52 13.7
Sequential - - - 78.9 64 12.7 1002 34.8 1.83 9.05
IterativeSamp - - - 85.4 64 18.7 1596 24.0 1.34 1.26
FALCONEye-Pro 4.3 1.7K 5.8 22.5 64 9.8 245.4 44.7 2.50 24.9
FALCONEye-Flash 3.0 1.7K 5.8 17.4 64 9.8 187.9 41.8 2.24 22.7
Baselines-.
A widely used baseline in recent long VQA benchmarks [chandrasegaran2024hourvideo] is the Socratic method [zeng2022socratic]. In this baseline, a VLM generates captions $\mathcal{C}=\{c_{1},\ldots,c_{n}\}$ from short video clips $\{v_{1},\ldots,v_{n}\}\in\mathcal{V}$ and a LLM generates $\mathcal{A}$ from $Q$ and $\mathcal{C}$. In addition to this baseline, we introduce three alternative exploration baselines (Fig. 7 of the supp. material) that rely solely on a VLM, guided by its answer confidence computed as described in Sec 3.
@@
As discussed in the related work, two relevant and recent video agents are VideoAgent [wang2024videoagent] and VideoTree [wang2025videotree]. We use VideoAgent as our reference meta‑architecture to compare with since it couples CLIP and a 7B‑parameter VLM, both of which fit comfortably on a single RTX3090. In contrast, VideoTree depends on EVA‑CLIP 8B—too large to co‑load with a VLM‑7B on that GPU—and its public codebase lacks the VLM captioning module, preventing a fair, fully reproducible baseline.
@@
We evaluate our meta-architecture using Qwen2.5-VL [Qwen2.5-VL] as the VLM and GPT4o-mini [openai2023chatgpt] as the LLM. As detailed in Sec. 6, Qwen2.5-VL outperforms other small-size VLMs on FALCON-Bench OQs in short clips near the answer, and is well calibrated. Its ViT supports images of any resolution, dynamically adjusting visual tokens, enabling an accurate zoom-in effect as clip lengths decrease. GPT4o-mini provides a balanced trade-off between cost and reasoning power (Sec. 6). Prompts used for LLM queries are detailed in Sec. 9.2. Inspired by the Socratic approach [chandrasegaran2024hourvideo], first-level clips are 1min length, but captions are now shorter formed by up to 64 tokens (instead of 768). Within these clips, second-order (20s) and third-order (5s) clips are explored. The zoom-in effect, validated in supplementary Sec. 11, adjusts from 30 frames ($672\times 1204$) in first-level clips to 10 frames ($824\times 1462$) in third-order clips. In OQs, the LLM determines the final answer based on completion, confidence ($0.8$ threshold), captions, and temporal localization (as explained in Sec. 3). In MCQs, this evaluation is replaced by a simple confidence-based check with $0.9$ threshold since all possible answers are predefined (A, B, C, or D) and inherently valid, making confidence the only selection criterion. Lastly, we evaluate two versions of FALCONEye: FALCONEye-Pro, the top-performing configuration, which evaluates up to 45 candidate clips, and FALCONEye-Flash, a cost-efficient variant that limits evaluation to 10 candidates.